# AutoAct: Automatic Agent Learning from Scratch via Self-Planning

### 10 Jan 2024 v1

@@ Abstract
Language agents have achieved considerable performance on various complex tasks. Despite the incessant exploration in this field, existing language agent systems still struggle with costly, non-reproducible data reliance and face the challenge of compelling a single model for multiple functions. To this end, we introduce AutoAct, an automatic agent learning framework that does not rely on large-scale annotated data and synthetic trajectories from closed-source models (e.g., GPT-4). Given limited data with a tool library, AutoAct first automatically synthesizes planning trajectories without any assistance from humans or strong closed-source models. Then, AutoAct leverages a division-of-labor strategy to automatically differentiate based on the target task information and synthesized trajectories, producing a sub-agent group to complete the task. We conduct comprehensive experiments with different LLMs, which demonstrates that AutoAct yields better or parallel performance compared to various strong baselines. We even notice that AutoAct, when using the Llama-2-13b model, can achieve performance comparable to that of the GPT-3.5-Turbo agent [^1] Code will be available at https://github.com/zjunlp/AutoAct..
@@ 1 Introduction
Language agents Wang et al. (2023b); Xi et al. (2023), which leverage the powerful reasoning capabilities Qiao et al. (2023b); Zhang et al. (2023b) of Large Language Models (LLMs) to generate executable actions for observing the external world, have emerged as essential components of AI systems designed to address intricate interactive tasks Torantulino (2023); Osika (2023); Nakajima (2023); Tang et al. (2023). The process of endowing LLMs with such interactive capabilities is referred to as Agent Learning wherein planning plays a pivotal role, which is responsible for decomposing complex tasks Wei et al. (2022); Zhou et al. (2023a); Yao et al. (2023), invoking external tools Shen et al. (2023); Lu et al. (2023); Qin et al. (2023), reflecting on past mistakes Shinn et al. (2023); Madaan et al. (2023), and aggregating information from various sources to reach the final targets. There have been a lot of works Yao et al. (2023); Wang et al. (2023a); Shen et al. (2023); Chen et al. (2023c) that directly prompt closed-source off-the-shelf LLMs to plan on particular tasks. Despite their convenience and flexibility, closed-source LLMs inevitably suffer from unresolved issues, as their accessibility often comes at a steep price and their black-box nature makes the result reproduction difficult. In light of this, some recent endeavors have shifted their focus towards imbuing open-source models with planning capabilities through fine-tuning Chen et al. (2023a); Zeng et al. (2023); Yin et al. (2023).
@@
However, despite the existing fine-tuning-based method’s achievements, they are not without their limitations. On the one hand, training open-source models necessitates a substantial amount of annotated task data and still relies on closed-source models to synthesize planning trajectories. However, fulfilling these requirements in many private scenarios, such as personal privacy and company business, often proves to be rocky. On the other hand, from the perspective of agent framework design, fine-tuning-based methods compel one single language agent to learn all planning abilities from the data, placing even greater pressure on them. These contradict the famous Simon’s principle of bounded rationality Mintrom (2015), which states that “precise social division-of-labor and clear individual tasks can compensate for the limited ability of individuals to process and utilize information”.
To this end, we introduce AutoAct, an automatic agent learning framework, which does not rely on large-scale annotated data and synthetic trajectories from closed-source models while incorporating explicit individual tasks and precise division-of-labor (see Figure 1). AutoAct starts with a Meta-Agent to obtain an augmented database through self-instruct Wang et al. (2023c) based on a limited set of user-provided data examples. Then, armed with a prepared tool library, the Meta-Agent can automatically synthesize planning trajectories without any assistance from humans or strong closed-source models. Finally, we propose the division-of-labor strategy which resembles cell differentiation based on the self-synthesized trajectories (genes), where the Meta-Agent acts as a stem cell Colman (2008) and differentiates into three sub-agents with distinct functions: task decomposition, tool invocation, and self-reflection, respectively. Our differentiation process is essentially a parameter-efficient training process on the self-synthesized trajectories which is lightweight and low consumption. We list the differences between AutoAct and prior works in Table 1.
Experiments on complex question-answering tasks with different LLMs demonstrate that AutoAct yields better or parallel performance compared to various strong baselines. We summarize our main contributions as follows:
@@
  • We conduct comprehensive experiments with different LLMs, which demonstrates that AutoAct yields better or parallel performance compared to various strong baselines. We even notice that when using the Llama-2-13b model, AutoAct can achieve performance comparable to that of the GPT-3.5-Turbo agent OpenAI (2022).
  • Extensive empirical analysis demonstrates the effectiveness of our appropriate division-of-labor strategy and the trajectory quality generated by AutoAct significantly outperforms that of other methods from multiple aspects.
@@ 2 AutoAct
Figure 2: The overview of our proposed framework AutoAct. We initiate with self-instruct to extend the task database from scratch. Then self-planning is applied to conduct automatic agent learning, including automatic tool selection, trajectories synthesis, self-differentiation and group planning. Our self-differentiation is a parameter-efficient fine-tuning process to achieve lightweight and low consumption.
@@ 2.1 Critical Components of AutoAct
The Meta-Agent stands at the central position of our AutoAct framework. It is responsible for all the preparatory work before self-differentiation and serves as the backbone model for the self-differentiation process. Equipped with a pre-prepared tool library, the Meta-Agent can differentiate into an agent group capable of collaborating to accomplish the target task, based on the limited target task information available. The Meta-Agent used in our approach can be initialized with any open-source model. In this paper, we primarily utilize the Llama-2 model series.
@@
Due to our agent learning from scratch, the task information at hand is quite limited, primarily encompassing three aspects: task name $\mathcal{M}$, task description $\mathcal{P}$, task data examples $\mathcal{C}$. Concretely, $\mathcal{P}$ represents a detailed description of the task’s characteristics, properties, and other relevant information. $\mathcal{C}=\{q_{i},a_{i}\}_{i=1}^{|\mathcal{C}|}$ indicates $|\mathcal{C}|$ question-answer example pairs of the task, where $|\mathcal{C}|$ is a much small value which users can effortlessly provide. For a more in-depth view of task information, please refer to Appendix B. Note that the task information serves as the only user-provided knowledge of the task for AutoAct to conduct automatic agent learning.
@@
To facilitate our agents in automatic task planning, we provide a comprehensive tool library at their disposal. The tool library can be denoted as $\mathcal{T}=\{m_{i},d_{i},u_{i}\}_{i=1}^{|\mathcal{T}|}$, where $m$ represents the name of each tool, $d$ defines the functionality of each tool, $u$ details the usage instruction of each tool, and $|\mathcal{T}|$ stands for the tools amount of the library. In our automatic procedure, the Meta-Agent has the autonomy to select appropriate tools from the tool library based on the task information. Users also have the option to expand the tool library according to their specific needs, allowing for more flexible utilization.
@@ 2.3 Automatic Agent Learning via Self-Planning
where $\tau^{r}$ and $\alpha^{r}$ represent the thought and action of the reflection process respectively, and $\mathcal{H}$ is the planning history after finishing the answer. The trajectories can be reorganized based on the responsibilities above and fed to the Meta-Agent for self-differentiation. Our differentiation is a parameter-efficient fine-tuning process to achieve lightweight and low consumption. Particularly, for each sub-agent, we train a specific LoRA Hu et al. (2022).
@@
After obtaining the task-specific sub-agents, any new question is processed through group planning among the sub-agents to achieve the desired outcome. Once the tool name $\alpha_{t}^{m}$ generated by the Plan-Agent is triggered at time $t$, the Tool-Agent is roused to determine the parameters $\alpha_{t}^{p}$ transferred to the specific tool. The return result of the tool is treated as the observation $o_{t}$ and handed to the Plan-Agent. After the collaboration between the Plan-Agent and Tool-Agent finishes with a prediction, the Reflect-Agent comes on the stage to reflect on the history and provide a reflection result contained in the reflection action $\alpha^{r}$. If the reflection result indicates that the prediction is correct, the whole planning process concludes. Otherwise, the Plan-Agent and Tool-Agent will continue the planning based on the reflection information.
@@ 3 Experimental Setup
We choose the open-source Llama-2 models Touvron et al. (2023) as the backbones of our Meta-Agent and sub-agents. The compared baselines are as follows: 1) CoT Wei et al. (2022), the naive Chain-of-Thought reasoning method. 2) ReAct Yao et al. (2023), a well-known single-agent framework based on few-shot learning that performs planning and action iteratively. 3) Reflexion Shinn et al. (2023), a single-agent framework to reinforce language agents through linguistic feedback. 4) Chameleon Lu et al. (2023), another few-shot single-agent framework that performs planning before action. 5) FireAct Chen et al. (2023a), a single-agent framework with fine-tuning on diverse kinds of trajectories generated by GPT-4 OpenAI (2023). 6) BOLAA Liu et al. (2023), a multi-agent architecture that customizes different agents through prompts. 7) GPT-3.5-Turbo OpenAI (2022), a strong closed-source baseline. To ensure fairness, we maintain an equal training trajectory volume of 200 for FireAct and AutoAct. As Reflexion provides answer correctness labels during reflection but other methods including AutoAct do not, we test all the other methods twice and choose the correct one for evaluation. For all the few-shot prompt-based baselines, we uniformly provide two examples in the prompt.
@@
We fine-tune all our models with LoRA Hu et al. (2022) in the format proposed in Alpaca Taori et al. (2023). Our fine-tuning framework leverages FastChat Zheng et al. (2023) using DeepSpeed Rasley et al. (2020). We set the learning rate of 1e-4 and the sequence length of 4096 for all the model scales. The training epoch for the 13b and 7b models is 5 and for the 70b model is 3. The batch size is 4 for the 13b and 7b models, and 1 for the 70b model. We use the AdamW optimizer Loshchilov and Hutter (2019) with a cosine learning scheduler. We detail the hyper-parameters for training in Appendix A.
@@ 4 Results
As shown in Table 2, the 13b and 70b models consistently outperform various prompt-based baselines. The 70b model even surpasses the agent performance of GPT-3.5-Turbo, achieving a rise of $\uparrow$3.77on HotpotQA and $\uparrow$6.39on ScienceQA. The performance of the 7b model is comparable to other methods to some extent. However, it fails to deliver the same impressive results due to its limited self-instruct and self-planning capabilities. Therefore, whether in a single-agent or multi-agent architecture, prompt-based methods relying on few-shot prompting fail to precisely customize the behavior of the agent, which is also supported by the fact that FireAct widely outperforms ReAct and BOLAA in the context of iterative planning. In addition, our investigation reveals a visible disparity in open-source models between the performance of many prompt-based planning baselines (relying on various external tools) and CoT (relying on the models’ intrinsic reasoning abilities). This discrepancy underscores the formidable challenge of unlocking planning capabilities by prompting.
@@
Further focusing on FireAct in Table 2, despite the assistance of GPT-4, FireAct’s approach of assigning the entire planning task to a single model proves to be burdensome. As a result, its performance on ScienceQA even falls short compared to the prompt-based global planning method, Chameleon. AutoAct employs self-differentiation to decouple the planning process and reaches a clear division-of-labor among sub-agents for group planning, resulting in a significant improvement than FireAct, with an enhancement of $\uparrow$5.77on HotpotQA and $\uparrow$6.67on ScienceQA with 70b model. Additionally, AutoAct achieves self-planning without relying on closed-source models and large-scale labeled datasets, which paves the way for automatic agent learning with open-source models from scratch. In ablation study (§4) and human evaluation (§5), we will further validate that the quality of trajectories synthesized by AutoAct is not inferior to FireAct trained on trajectories synthesized using GPT-4.
@@
Table 3 presents the performance of AutoAct on the 70b model after removing certain key processes. It can be observed that the least impactful removal is the - reflect. We investigate that in the zero-shot scenario, the model tends to be over-confident in its answers. It typically only recognizes its errors when there are obvious formatting mistakes or significant repetitions in the planning process. Consistent with previous findings, the removal of the - multi agents leads to a noticeable decrease in performance. A more exciting discovery is that the results of - multi are comparable to those of FireAct. This indirectly suggests that the trajectory quality generated by the 70b model is no worse than that of GPT-4. As expected, the performance significantly deteriorates after - fine-tuning, which once again confirms the previous conclusion. To demonstrate the necessity of filtering out planning error data, we specifically remove the filtering process (- filtering) to examine the performance of AutoAct. The results indicate that the damage caused by training on unfiltered data is even greater than that of - fine-tuning.
@@ 6 Related Work
The rise of LLMs has positioned them as the most promising key to unlocking the door to Artificial General Intelligence (AGI), providing robust support for the development of LLM-centered AI agents Wang et al. (2023b); Xi et al. (2023). Related works focus primarily on agent planning Yao et al. (2023); Song et al. (2022); Chen et al. (2023a), external tools harnessing Patil et al. (2023); Qiao et al. (2023a); Qin et al. (2023), collective intelligence among multi-agents Liang et al. (2023); Liu et al. (2023); Chen et al. (2023b), human and social property inside agents Zhang et al. (2023a); Park et al. (2023); Mao et al. (2023), etc. However, despite their success, existing methods still face two major troubles. Firstly, most agents heavily rely on prompts for customization, which makes it difficult to precisely tailor the behavior of the agent, resulting in unexpected performance at times. Secondly, each agent is compelled to master all skills, making it challenging for the agent to achieve expertise in every domain. To mitigate the above issues, our approach leverages a proper division-of-labor strategy and fine-tuning each sub-agent to equip different agents with distinct duties. These agents collaborate to accomplish tasks orderly and effectively.

### 17 Jan 2024 v2

@@ Abstract
Language agents have achieved considerable performance on various complex tasks. Despite the incessant exploration in this field, existing language agent systems still struggle with costly, non-reproducible data reliance and face the challenge of compelling a single model for multiple functions. To this end, we introduce AutoAct, an automatic agent learning framework that does not rely on large-scale annotated data and synthetic trajectories from closed-source models (e.g., GPT-4). Given limited data with a tool library, AutoAct first automatically synthesizes planning trajectories without any assistance from humans or strong closed-source models. Then, AutoAct leverages a division-of-labor strategy to automatically differentiate based on the target task information and synthesized trajectories, producing a sub-agent group to complete the task. We conduct comprehensive experiments with different LLMs, which demonstrates that AutoAct yields better or parallel performance compared to various strong baselines. We even notice that AutoAct, when using the Llama-2-13b model, can achieve performance comparable to that of the zero-shot GPT-3.5-Turbo agent [^1] Code will be available at https://github.com/zjunlp/AutoAct..
@@ 1 Introduction
Language agents Wang et al. (2023b); Xi et al. (2023); Xie et al. (2023), which leverage the powerful reasoning capabilities Qiao et al. (2023b); Zhang et al. (2023b) of Large Language Models (LLMs) to generate executable actions for observing the external world, have emerged as essential components of AI systems designed to address intricate interactive tasks Torantulino (2023); Osika (2023); Nakajima (2023); Tang et al. (2023). The process of endowing LLMs with such interactive capabilities is referred to as Agent Learning wherein planning plays a pivotal role, which is responsible for decomposing complex tasks Wei et al. (2022); Zhou et al. (2023a); Yao et al. (2023), invoking external tools Shen et al. (2023); Lu et al. (2023); Qin et al. (2023), reflecting on past mistakes Shinn et al. (2023); Madaan et al. (2023), and aggregating information from various sources to reach the final targets. There have been a lot of works Yao et al. (2023); Wang et al. (2023a); Shen et al. (2023); Chen et al. (2023c) that directly prompt closed-source off-the-shelf LLMs to plan on particular tasks. Despite their convenience and flexibility, closed-source LLMs inevitably suffer from unresolved issues, as their accessibility often comes at a steep price and their black-box nature makes the result reproduction difficult. In light of this, some recent endeavors have shifted their focus towards imbuing open-source models with planning capabilities through fine-tuning Chen et al. (2023a); Zeng et al. (2023); Yin et al. (2023).
@@
However, despite the existing fine-tuning-based method’s achievements, they are not without their limitations. On the one hand, training open-source models necessitates a substantial amount of annotated task data and still relies on closed-source models to synthesize planning trajectories. However, fulfilling these requirements in many real-world scenarios, such as private personal bot or sensitive company business, often proves to be rocky. On the other hand, from the perspective of agent framework design, fine-tuning-based methods compel one single language agent to learn all planning abilities, placing even greater pressure on them. These contradict the Simon’s principle of bounded rationality Mintrom (2015), which states that “precise social division-of-labor and clear individual tasks can compensate for the limited ability of individuals to process and utilize information”.
To this end, we introduce AutoAct, an automatic agent learning framework, which does not rely on large-scale annotated data and synthetic trajectories from closed-source models while incorporating explicit individual tasks with precise division-of-labor (see Figure 1). Given a limited set of user-provided data examples, AutoAct starts with a Meta-Agent to obtain an augmented database through self-instruct Wang et al. (2023c). Then, armed with a prepared tool library, the Meta-Agent can automatically synthesize planning trajectories without any assistance from humans or strong closed-source models. Finally, we propose the division-of-labor strategy which resembles cell differentiation based on the self-synthesized trajectories (genes), where the Meta-Agent acts as a stem cell Colman (2008) and differentiates into three sub-agents with distinct functions: task decomposition, tool invocation, and self-reflection, respectively. Our differentiation process is essentially a parameter-efficient training process on the self-synthesized trajectories with low consumption resources. We list the differences between AutoAct and prior works in Table 1.
Experiments on complex question-answering tasks with different LLMs demonstrate that the proposed AutoAct yields better or parallel performance compared to various strong baselines. We summarize our main contributions as follows:
@@
  • We conduct comprehensive experiments with different LLMs, which demonstrates that AutoAct yields better or parallel performance compared to various strong baselines. We even notice that when using the Llama-2-13b model, AutoAct can achieve performance comparable to that of the zero-shot GPT-3.5-Turbo agent OpenAI (2022).
  • Extensive empirical analysis demonstrates the effectiveness of our appropriate division-of-labor strategy and the trajectory quality generated by AutoAct outperforms that of other methods from multiple aspects.
@@ 2 AutoAct
Figure 2: The overview of our proposed framework AutoAct. We initiate with self-instruct to extend the task database from scratch. Then self-planning is applied to conduct automatic agent learning, including automatic tool selection, trajectories synthesis, self-differentiation and group planning. Our self-differentiation is a parameter-efficient fine-tuning process to achieve resource-efficient learning.
@@ 2.1 Critical Components of AutoAct
The Meta-Agent stands at the central position of our AutoAct framework. It is responsible for all the preparatory work before self-differentiation and serves as the backbone model for the self-differentiation process. Given limited target task information and a pre-prepared tool library, the Meta-Agent can differentiate into an agent group capable of collaborating to accomplish the target task. In AutoAct, the Meta-Agent can be initialized with any open-source model.
@@
In this paper, we mainly focus on agent learning from scratch, which means the task information at hand is quite limited, primarily encompassing three aspects: task name $\mathcal{M}$, task description $\mathcal{P}$, task data examples $\mathcal{C}$. Concretely, $\mathcal{P}$ represents a detailed description of the task’s characteristics, properties, and other relevant information. $\mathcal{C}=\{q_{i},a_{i}\}_{i=1}^{|\mathcal{C}|}$ indicates $|\mathcal{C}|$ question-answer example pairs of the task, where $|\mathcal{C}|$ is very small which users can effortlessly provide (e.g., a few demonstrations). For a more in-depth view of task information, please refer to Appendix B. Note that the task information serves as the only user-provided knowledge of the task for AutoAct to conduct automatic agent learning.
@@
To facilitate our agents in automatic task planning, we provide a comprehensive tool library at their disposal. The tool library can be denoted as $\mathcal{T}=\{m_{i},d_{i},u_{i}\}_{i=1}^{|\mathcal{T}|}$, where $m$ represents the name of each tool, $d$ defines the functionality of each tool, $u$ details the usage instruction of each tool, and $|\mathcal{T}|$ stands for the tools amount of the library. In our automatic procedure, the Meta-Agent has the autonomy to select appropriate tools from the tool library based on the task information. Users also have the option to expand the tool library according to their specific needs, allowing for more flexible utilization. We list part of our tool library in Appendix C.
@@ 2.3 Automatic Agent Learning via Self-Planning
where $\tau^{r}$ and $\alpha^{r}$ represent the thought and action of the reflection process respectively, and $\mathcal{H}$ is the planning history after finishing the answer. The trajectories can be reorganized based on the responsibilities above and fed to the Meta-Agent for self-differentiation. Our differentiation is a parameter-efficient fine-tuning process to achieve resource-efficient learning. Particularly, for each sub-agent, we train a specific LoRA Hu et al. (2022).
@@
After obtaining the task-specific sub-agents, any new question is processed through group planning among the sub-agents to achieve the desired outcome. Once the tool name $\alpha_{t}^{m}$ generated by the Plan-Agent is triggered at time $t$, the Tool-Agent is roused to determine the parameters $\alpha_{t}^{p}$ transferred to the specific tool. The return result of the tool is treated as the observation $o_{t}$ and handed to the Plan-Agent. After the collaboration between the Plan-Agent and Tool-Agent finishes with a prediction, the Reflect-Agent comes on the stage to reflect on the history and provide a reflection result contained in the reflection action $\alpha^{r}$. If the reflection result indicates that the prediction is correct, the whole planning process concludes. Otherwise, the Plan-Agent and Tool-Agent will continue the planning based on the reflection information. The specific sequence of the group planning process can be referred to the trajectory example on the right side of Figure 2.
@@ 3 Experimental Setup
We choose the open-source Llama-2 models Touvron et al. (2023) as the backbones of our Meta-Agent and sub-agents. The compared baselines are as follows: 1) CoT Wei et al. (2022), the naive Chain-of-Thought reasoning method. 2) ReAct Yao et al. (2023), a well-known single-agent framework based on few-shot learning that performs planning and action iteratively. 3) Chameleon Lu et al. (2023), another few-shot single-agent framework that performs planning before action. 4) Reflexion Shinn et al. (2023), a single-agent framework to reinforce language agents through linguistic feedback. 5) BOLAA Liu et al. (2023), a multi-agent framework that customizes different agents through prompts. 6) FireAct Chen et al. (2023a), a single-agent framework with fine-tuning on diverse kinds of trajectories generated by GPT-4 OpenAI (2023). 7) GPT-3.5-Turbo OpenAI (2022). To ensure fairness, we maintain an equal training trajectory volume of 200 for FireAct and AutoAct (200 synthesized data). As Reflexion provides answer correctness labels during reflection but other methods including AutoAct do not, we test all the other methods twice and choose the correct one for evaluation. For all the prompt-based baselines, we uniformly provide two examples in the prompt.
@@
We fine-tune all our models with LoRA Hu et al. (2022) in the format proposed in Alpaca Taori et al. (2023). Our fine-tuning framework leverages FastChat Zheng et al. (2023) using DeepSpeed Rasley et al. (2020). We set the learning rate of 1e-4 and the sequence length of 4096 for all the model scales. The training epoch for the 13b and 7b models is 5 and for the 70b model is 3. The batch size is 4 for the 13b and 7b models and 1 for the 70b model. We use the AdamW optimizer Loshchilov and Hutter (2019) with a cosine learning scheduler. We detail the hyper-parameters for training in Appendix A.
@@ 4 Results
As shown in Table 2, the 13b and 70b models consistently outperform various prompt-based baselines. The 70b model even surpasses the agent performance of GPT-3.5-Turbo, achieving a rise of $\uparrow$3.77on HotpotQA and $\uparrow$6.39on ScienceQA. The performance of the 7b model is comparable to other methods to some extent. However, it fails to deliver the same impressive results due to its limited self-instruct and self-planning capabilities. Therefore, whether in a single-agent or multi-agent architecture, prompt-based methods relying on few-shot demonstrations fail to precisely customize the behavior of the agent, which is also supported by the fact that FireAct widely outperforms ReAct and BOLAA in the context of iterative planning. In addition, our investigation reveals a visible disparity in open-source models between the performance of many prompt-based planning baselines (relying on various external tools) and CoT (relying on the models’ intrinsic reasoning abilities). This discrepancy underscores the formidable challenge of unlocking planning capabilities by prompting.
@@
Further focusing on FireAct in Table 2, despite the assistance of GPT-4, FireAct’s approach of assigning the entire planning task to a single model proves to be burdensome. As a result, its performance on ScienceQA even falls short compared to the prompt-based global planning method, Chameleon. AutoAct employs self-differentiation to decouple the planning process and reaches a clear division-of-labor among sub-agents for group planning, resulting in a improvement than FireAct, with an enhancement of $\uparrow$5.77on HotpotQA and $\uparrow$6.67on ScienceQA with 70b model. Additionally, AutoAct achieves self-planning without relying on closed-source models and large-scale labeled datasets, which paves the way for automatic agent learning with open-source models from scratch. In ablation study (§4) and human evaluation (§5), we will further validate that the quality of trajectories synthesized by AutoAct is not inferior to FireAct trained on trajectories synthesized using GPT-4.
@@
Table 3 presents the performance of AutoAct on the 70b model after removing certain key processes. It can be observed that the least impactful removal is the - reflect. We investigate that in the zero-shot scenario, the model tends to be over-confident in its answers. It typically only recognizes its errors when there are obvious formatting mistakes or significant repetitions in the planning process. Consistent with previous findings, the removal of the - multi agents leads to a noticeable decrease in performance. A more exciting discovery is that the results of - multi are comparable to those of FireAct. This indirectly suggests that the trajectory quality generated by the 70b model may be no worse than that of GPT-4. As expected, the performance deteriorates after - fine-tuning, which once again confirms the previous conclusion. To demonstrate the necessity of filtering out planning error data, we specifically remove the filtering process (- filtering) to examine the performance of AutoAct. The results indicate that the damage caused by training on unfiltered data is even greater than that of - fine-tuning.
@@ 6 Related Work
The rise of LLMs has positioned them as the most promising key to unlocking the door to Artificial General Intelligence (AGI), providing robust support for the development of LLM-centered AI agents Wang et al. (2023b); Xi et al. (2023); Wang et al. (2023d, e). Related works focus primarily on agent planning Yao et al. (2023); Song et al. (2022); Chen et al. (2023a), external tools harnessing Patil et al. (2023); Qiao et al. (2023a); Qin et al. (2023), collective intelligence among multi-agents Liang et al. (2023); Liu et al. (2023); Chen et al. (2023b), human and social property inside agents Zhang et al. (2023a); Park et al. (2023); Mao et al. (2023), etc. However, despite their success, existing methods still face two major troubles. Firstly, most agents heavily rely on prompts for customization, which makes it difficult to precisely tailor the behavior of the agent, resulting in unexpected performance at times. Secondly, each agent is compelled to master all skills, making it challenging for the agent to achieve expertise in every domain. In response, our approach leverages a proper division-of-labor strategy and fine-tuning each sub-agent to equip different agents with distinct duties. These agents collaborate to accomplish tasks orderly and effectively.